### v1

# A Survey of Deep Learning-based Radiology Report Generation Using Multimodal Data

Abstract
Automatic radiology report generation can alleviate the workload for physicians and minimize regional disparities in medical resources, therefore becoming an important topic in the medical image analysis field. It is a challenging task, as the computational model needs to mimic physicians to obtain information from multi-modal input data (i.e., medical images, clinical information, medical knowledge, etc.), and produce comprehensive and accurate reports. Recently, numerous works emerged to address this issue using deep learning-based methods, such as transformers, contrastive learning, and knowledge-base construction. This survey summarizes the key techniques developed in the most recent works and proposes a general workflow for deep learning-based report generation with five main components, including multi-modality data acquisition, data preparation, feature learning, feature fusion/interaction, and report generation. The state-of-the-art methods for each of these components are highlighted. Additionally, training strategies, public datasets, evaluation methods, current challenges, and future directions in this field are summarized. We have also conducted a quantitative comparison between different methods under the same experimental setting. This is the most up-to-date survey that focuses on multi-modality inputs and data fusion for radiology report generation. The aim is to provide comprehensive and rich information for researchers interested in automatic clinical report generation and medical image analysis, especially when using multimodal inputs, and assist them in developing new algorithms to advance the field.
Keywords Report generation $\cdot$ Deep learning $\cdot$ Multimodal $\cdot$ Medical image analysis
1 Introduction
Medical images can offer detailed insights into bodies and help physicians screen, diagnose and monitor medical conditions without requiring invasive techniques (Beddiar et al., 2023; Liao et al., 2023). Radiologists summarize the information extracted from medical imaging into radiological reports for clinical decision-making. The manual generation of reports is however labour-intensive, time-consuming, and requires extensive expertise (Beddiar et al., 2023). Topol (2019) points out that the demand for medical image explanation greatly surpasses the current capacity of physicians in the United States. During an epidemic and with ageing populations, the situation can get worse. During the Covid-19 pandemic, for instance, in the UK, each radiologist was estimated to report as many as 100 images each day (Statistics, 2020). This makes it challenging for radiologists to provide high-quality reports within the scheduled time. The current demand extends patient waiting time and increases the risk of disease transmission (Beddiar et al., 2023) and compromises patient care. The development of automatic report generation techniques can help alleviate this problem.
Automatic high-quality report generation is challenging. It is intrinsically a multi-modality problem (Tu et al., 2024; Yan et al., 2023). In routine clinical practice, to produce clear, correct, concise, complete, consistent, and coherent reports, radiologists need to combine information from images with information from other modality data, such as clinical history and related clinical measures. Previously developed techniques mostly considered images as input, while for the past three years, multi-modality deep learning developed very rapidly. An increasing number of research papers endeavored to emulate physicians by leveraging multi-modal data for the generation of diagnostic reports, as shown in Figure 1.
Figure 1: The distributions of reviewed papers using image data and multi-modality data as inputs per year from 2021 to 2024. The percentage denotes the input’s prevalence among articles published within the year.
Most of the previous surveys on this topic (Kaur et al., 2022; Beddiar et al., 2023; Liao et al., 2023; Shamshad et al., 2023; Liu et al., 2023a) did not include non-imaging inputs. Messina et al. (2022) considered non-imaging inputs, but only involved 6 papers. Totally, the previous surveys included 40 to 66 papers for report generation, primarily focusing on articles published before 2022. This survey differs from previous ones in three main contributions: (1) we analyse an additional 22 papers that utilize non-image inputs, and focuses on the acquisition, analysis, and integration of multi-modal inputs. To the best of our knowledge, it is the first review to investigate state-of-the-art multi-modal data processing techniques for report generation; (2) we examine 89 papers published from 2021 to 2024 to provide a comprehensive study on novel techniques in automatic report generation; and (3) we propose a general workflow for report generation with a taxonomy of approaches employed, and we summarize training strategies, public datasets, and mainstream evaluation methods, as shown in Figure 2. The workflow includes 5 key components: inputs, data preparation, feature learning, feature fusion, and report generation. Table LABEL:tab:overall in Appendix A summarizes all works included in this survey.
Figure 2: The summarized workflow of automatic radiology report generation. The fundamental components and key techniques are included. The (x, y%) for each method represents the number and percentage of papers used that technique in our survey.
The remainder of the paper is organized as follows. Section 2 introduces the paper search and selection process. Section 3 first provides a workflow of deep learning-based report generation, then analyses the techniques in each component of the workflow, and finally introduces the overall training strategies. Next, in sections 4 and 5, we introduce popular public datasets and evaluation methods, including metrics and expert evaluation. Section 6 compares the model performance of several papers under the same experimental setting. Lastly, we discuss challenges and perspectives on this topic in section 7 and provide a conclusion in section 8.
2 Search and selection of articles
Three search engines (Google Scholar, PubMed, and Springer) and four queries were employed to collect articles. They included “radiology report generation", (medical OR medicine OR health OR radiology) AND (report OR description OR caption) AND generation, modal AND (medical OR medicine OR health OR radiology) AND (report OR description OR caption) AND generation, and “medical report generation". Following the searches, the titles and abstracts of each article were read briefly to identify those that met the selection criteria. If there was uncertainty, the article was included to ensure relevant studies were not omitted. The selection criteria were framed around three aspects. First, we included articles published in the years 2021, 2022, 2023, and 2024 due to the significant number of developments using multi-modal technology in recent years. We aim to focus on the latest algorithms not covered in previous surveys. Second, the studies must be original researches focused on the automatic generation of full-text natural language radiology reports and include quantitative evaluation results. Techniques generating short captions of one or two sentences are excluded due to the differing nature of long report and sentence generation. Third, papers published in journals, conferences, and conference workshop proceedings were included. Moreover, papers uploaded on the arXiv website in 2023 and 2024 with over 30 citations were also selected. A total of 144 papers were identified using three search engines. In addition, by tracing the ancestry and descendants of papers, we identified another 24 papers. After removing duplicates, 97 publications were retained. We thoroughly read these works and applied exclusion criteria. First, at least one of the generated languages should be English. Second, at least one of the input data should be images. Finally, 89 works were included in the following analysis.
3 Methods
Deep learning-based radiology report generation typically follows a standard workflow summarized in Figure 2. This section analyzes the techniques in the 89 works, based on the workflow identified. Overall, a basic radiology report generation framework consists of 5 steps: (1) multi-modality data acquisition (section 3.1); (2) data preparation (section 3.2); (3) feature learning (section 3.3); (4) feature fusion and interaction (section 3.4); and (5) report generation (Section 3.5). In addition, novel training strategies, including modifying loss functions, reinforcement learning, and curriculum learning, are described in Section 3.6.
Medical images, when analyzed with or without other types of data, are firstly prepared (step 2). Subsequently, they are input into feature extractors to perform feature learning (step 3), predominantly implemented using CNN or Transformer architectures, along with multiple enhancement modules (e.g., auxiliary task and contrastive learning). The feature extractors aim at extracting features relevant to report generation, and the enhancement modules are utilized to improve the expressiveness of the features. For certain approaches, a feature fusion and interaction module (step 4) is subsequently applied to align cross-modal data and mitigate the negative effects caused by differences between the visual and textual domains. After fusion and interaction, the features are conveyed back to the feature extractor or directly input into the generator to generate the report (step 5). The training strategy is implemented to improve learning effectiveness during training. Table LABEL:tab:overall in Appendix A presents detailed information for each paper across the five steps and training strategies.
3.1 Multimodality input data
The input data refers to the data received by the report generation system. During model training and inference, the data can vary; for example, both images and real reports are used as inputs during training, while only images are used during inference (Shetty et al., 2023). However, the model learns from the distribution and features of the training data. If the testing data changes, mostly, the model will struggle to generalize, resulting in decreased performance. Therefore, in this section, we mainly introduce the acquisition of input data that is consistent between the training and testing phases in the reviewed papers. The input data includes:
  • Image data includes X-ray, magnetic resonance imaging (MRI), computed tomography (CT), ultrasound, gastrointestinal endoscope image, retinal image, and dermoscopy image. Most papers we reviewed focus on generating medical reports for chest X-ray images (82 works). Other than chest diagnosis, retinal image is the second most prevalent image modality (6 works). The retinal image includes lots of categories, e.g., fundus fluorescein angiography and color fundus photography. Other works focus on chest CT (5 works), gastrointestinal endoscope image (2 works), spine MRI (1 work), dermoscopy image (1 work), and breast ultrasound images (1 work). Although some images are non-radiological, such as ophthalmic images, we still include their report-generation techniques in this survey.
  • Medical terminology refers to medical terms and expressions, which are from keyword labels of images (Huang et al., 2021a, b, 2022; Liu et al., 2023b), or self-built corpora (Liu et al., 2021b; Cao et al., 2022, 2023; Xue et al., 2024; Gu et al., 2024; Li et al., 2023c), storing common descriptions found in medical reports.
  • Medical knowledge base mostly records the connections between different organs and diseases, and is presented in a graph format. Graph is a fundamental data structure consisting of a set of nodes and edges, which can easily represent a set of subjects and their connections. The knowledge graph can primarily be obtained in two ways: 1) public datasets such as the RadGraph (Jain et al., 2021) used by two works (Yang et al., 2022; Li et al., 2023a); 2) self-designed knowledge bases according to authoritative medical standards (Huang et al., 2023; Xu et al., 2023) or disease labels (Jia et al., 2022). Yang et al. (2023) argued that manual graph construction is limited to diseases, further complicating the adaptation of these models for diverse datasets. To overcome it, they automatically constructed the medical knowledge based on real reports during training and leveraged the knowledge base during model inference.
  • Real texts report is mostly obtained by data retrieval (Liu et al., 2021b; Song et al., 2022; Yang et al., 2022; Li et al., 2023a; Liu et al., 2023b; Li et al., 2023c; Jin et al., 2024). For example, for each input image, Liu et al. (2021b) retrieved similar images from the training dataset and utilized the corresponding reports. This process mimics radiologists consulting previous medical case reports when drafting their own. In addition, Liu et al. (2023b) also obtained pre-defined sentences based on input terminologies.
  • Clinical information encompasses patient demographics (e.g., age and gender), clinical observations, and medical histories. It is included within the indication section of radiology reports (see Figure 4).
  • Questionnaires: report generation model can be trained in a visual question answering way (Tanwani et al., 2022; Pellegrini et al., 2023). The questionnaires are provided by public datasets, such as VQA-Rad (Lau et al., 2018) and Rad-ReStruct (Pellegrini et al., 2023).
3.2 Data preparation
Data preparation endeavors to enhance data quality and prepare it for model deployment, typically encompassing data cleansing, transformation, and organization. Conventional preparation methods include image resizing and cropping, text tokenizing, converting all tokens to lowercase, removing non-alphabetic tokens, and implementing data augmentation procedures. The methods utilized in each paper are outlined in Table LABEL:tab:overall in Appendix A, but it is worth noting that conventional preparation methods are so ubiquitous that some papers do not mention them. While the information is not recorded in the table, it does not mean the absence of a data preparation process.
Novel data preparation methods in the reviewed papers can be categorized into filtering (Ramesh et al., 2022) and grouping (Wang et al., 2022b). Ramesh et al. (2022) argued that writing a radiology report necessitates referencing historical information, which inevitably included descriptors such as ‘again’ and ‘decrease’. However, these terms cannot be inferred from a single image, therefore Ramesh et al. filtered such descriptions in the reports. This exclusion was found to facilitate the model’s learning process.
Grouping refers to organizing sentences from ground truth reports into distinct sections, typically relying on keywords from pre-defined knowledge graphs and filtering rules. Each section describes a specific anatomical structure. Grouping aims to enable the generation system to process various types of sentences differently. For instance, Wang et al. (2022b) employed different decoders to generate descriptions for different anatomical structures. Alongside the reviewed papers, a recently released public dataset named ImaGenome (Wu et al., 2021) also includes grouping results in their annotation files (see Section 4). The grouping result becomes more easily accessible.
3.3 Feature learning
3.3.1 Image-based feature learning
Previous research primarily utilized CNNs as architectures for extracting image features, however, recently, an increasing number of researchers have opted for the use of Transformers due to their improved performance. Simultaneously, numerous studies proposed novel modules to enhance the model capability. In this section, the model architecture is first introduced, and subsequently enhancement modules are described. These modules include auxiliary tasks, contrastive learning, and memory metrics. The architecture and modules utilized in each paper are outlined in Table LABEL:tab:overall in Appendix A.
CNN and Transformer encoder for feature extraction
The statistics of the architectures used as image feature extractor are shown in Figure 3. In total, forty-six studies extract image features purely based on CNN models. Thirty-four works firstly encode images by CNN and then utilizes the Transformer layers to modify the embeddings. Eight works utilize a pure Transformer architecture to extract image features. Figure 3 shows a clear trend of more studies adopting CNN augmented with Transformer for image feature extraction.
Figure 3: The statistics of the reviewed papers using different architectures to extract image features per year from 2021 to 2024. The percentage denotes the method’s prevalence among articles published within the year.
For CNN architecture, two works self-designed CNN models, while the other works built it based on different classical visual models, such as ResNet (He et al., 2016) (42 papers), DenseNet (Huang et al., 2017) (22 papers), VGG (Simonyan and Zisserman, 2014) (5 papers), Faster-RCNN (Ren et al., 2015) (5 papers), Inception-V3 (Szegedy et al., 2016) (1 paper), ResNeXt (Xie et al., 2017) (1 paper), EfficientNet (Tan and Le, 2019) (1 paper), and the Two-Stream Inflated 3D ConvNets (I3D) (Carreira and Zisserman, 2017) (1 paper). Pahwa et al. (2021) modified the HRNet (Sun et al., 2019a), a human pose estimation network. Other three works (Huang et al., 2021a, b, 2022) provided results based on different CNN structures.
To improve model performance, ten works modified the CNN structure by attention modules, which assigned varying degrees of importance (weights) to different parts of the input by learnable parameters, allowing the model to selectively focus on specific regions of an image. Traditional attention mechanisms can be classified into channel-wise (Du et al., 2022; Wang et al., 2022e; Gajbhiye et al., 2022; Pahwa et al., 2021) and spatial-wise (Pahwa et al., 2021; Jia et al., 2021), which allocate different weights to the various channels and spatial positions of the inputs respectively. In addition, Li et al. (2023b) and Wang et al. (2024a) utilized the idea of the class activation map (Zhou et al., 2016) to obtain weights. Yan et al. (2022) initially extracted image patch features, clustered them using an unsupervised method, and then weighted the cluster results. Experimental results show that attention mechanism allows models to pay more attention to the lesions than irrelevant background. With the rise of the Transformers (Vaswani et al., 2017), multi-head attention has become a potent method for information interaction. Wang et al. (2023c) first extracted regions of interest from the frontal view and then employed multi-head attention to fuse information between the frontal and lateral views with the regions. This approach introduced regions of interest into model training to improve model performance.
For Transformer architecture, most of them leverage a standard Transformer encoder, while Li et al. (2023d) argued that aligning images and text posed a challenge due to the continuous nature of images and the discrete nature of text; therefore, they improved the model performance by using a discrete variational autoencoder (Ramesh et al., 2021) to obtain discrete visual tokens. Other works improved the model performance by modifying the self-attention module. Three works (Wang et al., 2022e; Lin et al., 2023; Wang et al., 2023d) added high-order interactions among three inputs of the Transformer attention module. Two works (Miura et al., 2021; Wang et al., 2022e) were inspired by the memory-augmented attention (Cornia et al., 2020) and extended the keys and values with additional plain learnable vectors to record more information. Li et al. (2022b) introduced a learnable parameter in the attention operation. Wang et al. (2023d) modified the encoder by including additional input tokens. These tokens were named ‘expert tokens’ to emulate the “multi-expert joint diagnosis” methodology.
Auxiliary task for feature extraction
Auxiliary tasks aim to provide additional supervision signals to the feature extractor, enabling it to extract information relevant to report generation from images. These tasks mainly include classification (22 papers), graph construction (10 papers), embedding comparison (10 papers), and detection/segmentation (7 papers). Each of them is introduced in detail below.
Classification: The most common auxiliary task used in the reviewed papers is classification referring to assigning images to predefined categories. The predefined categories primarily include medical tags (Gajbhiye et al., 2022; Kaur and Mittal, 2022b; Wang et al., 2022e; Hou et al., 2021b; Du et al., 2022; You et al., 2022; Alfarghaly et al., 2021) and disease labels (Liu et al., 2021e; Zhou et al., 2021; Wang et al., 2022d; Yang et al., 2023; Zhang et al., 2023b; Wang et al., 2024a; Jin et al., 2024; Hou et al., 2021b). The medical tags are from standard medical vocabularies including hundreds of labels, such as anatomical structures and pathological signs. They are provided by manual annotations or auto annotation tools (e.g., NIH MTI web API  [^1] https://ii.nlm.nih.gov/MTI/index.shtml and RadGraph (Jain et al., 2021)). Disease labels are provided by auto annotation tools (e.g., CheXpert (Irvin et al., 2019) and CheXbert (Smit et al., 2020)). Compared to disease labels, medical tags offer a more comprehensive range of information. However, to the best of our knowledge, there is no literature that supports the superiority of medical tags over disease labels. Perhaps due to the extensive scope covered by medical tags, deep learning models face challenges in acquiring such rich knowledge. Zhou et al. (2021) incorporated 32 additional labels for lesion location, size, and shape (e.g., “upper/lower” and “patchy”) into the disease label set, observing a slight improvement in model performance.
Other notable categories used in the reviewed papers include matching status (Li et al., 2023a), local properties (Yang et al., 2021a), report cluster results (Li et al., 2022a), and fix answer categories (Tanwani et al., 2022). Li et al. (2023a) predicted the matching status of a given image-report pair. Yang et al. (2021a) devised localized property labels for breast ultrasound images, such as tumor morphology, to facilitate the identification of properties that are challenging to be discerned in low-resolution images. Li et al. (2022a) first conducted unsupervised clustering on the ground truth report, subsequently utilizing the resultant clusters as labels. This auxiliary task yielded a marked improvement in text generation performance. Tanwani et al. (2022) considered the report generation as a question-answer task, and the classifier was designed for fixed answer categories. In addition, for a detection auxiliary task, classifiers need to be applied to identify attributes of detected regions (e.g., ‘right lung’). This paragraph eschews such cases to circumvent redundancy.
Graph construction: Graph construction aims at introducing prior knowledge into the report generation process. The knowledge graph in this section differs from that in Section 3.1. Here, node features are extracted from images, and edges are defined as parameters in graph convolution networks. In contrast, the node and edge information in the input knowledge graphs is derived from non-image data. A classical method was proposed by Zhang et al. (2020) and yielded promising outcomes. A knowledge graph was constructed firstly based on insights provided by domain experts, where nodes represented major abnormalities and major organs, and bidirectional connections linked nodes that were related to each other. To initialize nodes features, a spatial attention module was introduced after the CNN backbone using 1$\times$1 convolution layers and softmax layers. The number of channels matched the number of nodes. The nodes’ initial embedding was derived as attention-weighted feature maps. Then graph convolution layers were employed to disseminate information throughout the graph, followed by two branches for classification and report generation. First, the classification branch was trained, and subsequently, parameters in both the CNN backbone and the graph convolution layers were frozen, only the report generation decoder was trained. Six works (Liu et al., 2021b, d; Cao et al., 2022; Wang et al., 2022c; Yan, 2022; Zhang et al., 2023b) utilized this method (Zhang et al., 2020). Wang et al. (2022c) expanded the graph (Zhang et al., 2020) by incorporating information from a radiology terms corpus named Radiology Lexicon (RadLex) [^2] http://www.radlex.org/ (Langlotz, 2006). As the number of graph nodes increased, the model performance initially improved, peaked at 40 nodes, and then declined, with a noticeable decrease at 60 nodes. Liu et al. (2021d) constructed a large graph based on the MIMIC-CXR dataset. The nodes represented frequent clinical abnormalities and the edges represented the co-occurrence situation of different abnormalities. In addition, Li et al. (2023b) used the disease prediction results to obtain the node features. The nodes were classification probabilities, with learnable edge weights.
Another graph reconstruction method aims at reconstructing triplets that are in the form of (entity1, relationship, entity2), such as (opacity, suggestive of, infection). Three works (Dalla Serra et al., 2022, 2023b; Li et al., 2022b) firstly predicted the triplets and then generated reports based on them. The experimental results show that using triplets alone for report generation is ineffective; combining them with features extracted from images is necessary for better results.
Embedding comparison: Embedding comparison refers to constraining the consistency of different features in intermediate layers, thereby guiding the learning process. The comparison in reviewed papers is mainly applied between features extracted from images and real reports (Najdenkoska et al., 2021, 2022; Zhou et al., 2021; Yang et al., 2021b; Chen et al., 2022; Wang et al., 2021, 2022d; Yang et al., 2023). Experimental results show that the supervision signals from real text enable extracted visual features carry richer semantic information, facilitating more effective translation into radiology reports. Four works (Wang et al., 2021; Zhou et al., 2021; Wang et al., 2022d; Yang et al., 2023) utilized a triple loss function to compel the image-text paired features to be closer to a latent space than the unpaired ones. Najdenkoska et al. (2021, 2022) inspired by the Auto-Encoding Variational Bayes (Kingma and Welling, 2013). They used real reports to obtain a latent space during training and generated reports based on this space. The image extractors were enabled to capture features from images that closely resemble those found in real reports. Other two works (Yang et al., 2021b; Chen et al., 2022) used the Term Frequency and Inverse Document Frequency (TF-IDF) to extract important information from real reports as supervision signals. TF-IDF is a statistical measure assessing a word’s importance by considering its frequency in a specific document and its rarity across the entire document set.
In addition, to produce reports for abnormalities not seen during training, Sun et al. (2022) initially linearly projected visual features to semantic features, and extracted semantic features of labels by the BioBert model (Lee et al., 2020). Consistent constrain was applied between two similarity: 1) the similarity between pairwise elements in the semantic features from visual features; and 2) the similarity between the semantic features from visual features and the semantic features from labels. Zhang et al. (2023b) integrated semi-supervised learning into report generation using two networks. They first applied different types of noise to an input image to create two variations, which were then fed into the two networks. An auxiliary loss function was employed to ensure consistency in the extracted visual features.
Detection/segmentation: Object detection locates and identifies objects or patterns within an image, focusing on determining their presence and position. Segmentation divides an image into meaningful regions by identifying and separating objects based on specific characteristics. Both processes enhance the model’s understanding of the image by object recognition and region extraction, and can improve the model’s interpretability by linking the detection/segmentation results with generated sentences. The detection/segmentation regions can be anatomical regions (Tanida et al., 2023; Dalla Serra et al., 2023b, a; Wang et al., 2023c; Han et al., 2021; Gu et al., 2024) and abnormal regions (Sun et al., 2022). No literature compares the impact of detection and segmentation tasks on report generation results. However, a publicly available dataset named Chest ImaGenome (see Section 4) offers detection annotations, making them easier to be acquired than segmentation annotations.
In addition, the outputs of auxiliary tasks can provide valuable information such as disease labels, therefore, inputting them into the following generation network is a common choice (Alfarghaly et al., 2021; Hou et al., 2021b; Singh et al., 2021; Yang et al., 2021a; You et al., 2021; Zhou et al., 2021; Du et al., 2022; Jia et al., 2022; Kaur and Mittal, 2022b; Sun et al., 2022; Wang et al., 2022a, e; Yan et al., 2022; You et al., 2022; Tanida et al., 2023; Li et al., 2022b; Jin et al., 2024). For example, Zhou et al. (2021) sent the semantic word embeddings of the predicted findings from the classifier to the report generation decoder.
Contrastive learning for feature extraction
Contrastive learning is a self-supervised learning method to improve the representational capacity of models, which allows models to minimize the distance among positive pairs and maximize it for negative ones. It can be used to train feature extractors (Wang et al., 2022e; Lin et al., 2023; Wu et al., 2022; Wang et al., 2023b). Lin et al. (2023) utilized a classical contrastive learning method named Momentum Contrast (He et al., 2020), where different views or augmented versions of the same image were considered as positive pairs. Another representative contrastive learning work is the Contrastive Language-Image Pre-training (CLIP) model (Radford et al., 2021). It connects textual and visual information by directly training on a vast dataset consisting of image-text pairs. Wang et al. (2022e) directly employed it for image feature extraction and Wu et al. (2022) applied the idea of CLIP to train the feature extractors on the training dataset. Wang et al. (2023b) however argued that previous works treated the entire report as input, overlooking the distinct information contained within individual sentences. This oversight could result in incorrect matching of image-text pairs. Therefore, they proposed phenotype-based contrastive learning. This method involved randomly initializing a set of vectors as phenotypes, allowing sentences and visual embeddings to interact with them, and finally conducting contrastive learning between the processed embeddings. The results outperformed previous contrastive learning methods in report generation.
Contrastive learning also can be part of the training loss (Tanwani et al., 2022; Wang et al., 2022a; Li et al., 2023a; Liu et al., 2023b), and can be applied between visual and textual features (i.e. image-text pairs) (Tanwani et al., 2022; Li et al., 2023a; Liu et al., 2023b), or be applied based on labels, treating samples with shared labels as positives and those without any common labels as negatives (Wang et al., 2022a). Their ablation study demonstrated a significant improvement in results compared to standard contrastive loss (Wang et al., 2022a).
Contrastive attention is another method to utilize contrastive learning. Liu et al. (2021c) designed a contrastive attention model to extract abnormal region features by comparing input samples with normal cases. Similar features shared between the input and normal cases were subtracted from the input image feature, and the remaining feature was then concatenated with the original feature. Song et al. (2022) argued that the contrastive technique (Liu et al., 2021c) did not consider historical information, therefore they proposed a module based on similarity retrieval technique to obtain similar images from the training dataset. The image features were processed by enlarging different features between inputs and the similar retrieved images.
Memory metric for feature extraction
Using a memory metric for image feature extraction assumes the presence of similar features in various medical images. Memory metrics are employed to record and transmit the similarity information during training (Chen et al., 2022; Yan, 2022). Typically, an n×n matrix is randomly initialized, where n represents the number of metric rows. Then, at each training step, the matrix is updated based on the visual features and the previous metrics. The memory metric used in this section is consistent with the metric used in the memory-driven transformer discussed in Section 3.5, with one being applied to images and the other to the generated reports.
3.3.2 Non-imaging-based feature learning
Most non-imaging data is presented in the form of text. Before fusion with image data, text data needs to be embedded. We first introduce a widely-used basic text embedding technique, the lookup table. This table assigns a unique index to each word or character, which is used to look up a pre-trained word vector or character vector. In addition to the basic method, additional feature extraction can be performed on these vectors to enhance their representation capabilities.
Transformer-based models such as BERT and its variants have emerged as mainstream methods for feature extraction across various textual data, such as terminology (Liu et al., 2021e; Cao et al., 2022, 2023; Liu et al., 2021b; Xue et al., 2024; Liu et al., 2023b; Li et al., 2023c), real text reports (Liu et al., 2021b, 2023b; Li et al., 2023c; Jin et al., 2024), knowledge graphs (Yang et al., 2022; Huang et al., 2023; Li et al., 2023a; Xu et al., 2023), and questionnaires (Tanwani et al., 2022; Pellegrini et al., 2023), and have achieved good results.
Several methods are designed for a specific type of input. Li et al. (2023c) used the TF-IDF to re-weight terminology embeddings. The re-weighted approach alleviated the issue of data imbalance, resulting in performance enhancement. Clinical information can be processed by a pre-trained feature extractor named BioSentVec (Zhou et al., 2021; Chen et al., 2019). For the knowledge base, Yang et al. (2022) used a knowledge graph embedding model named RotatE (Sun et al., 2019b) to obtain entity embeddings and relation embeddings from the RadGraph. Besides using the entire RadGraph as input, two works combined the real reports with RadGraph to extract case-related information from real reports and queried related information from the RadGraph (Yang et al., 2022; Li et al., 2023a). Alternatively, Li et al. (2023c) considered RadGraph as an annotation tool for extracting entities and positional information from real reports. Other two works (Xu et al., 2023; Huang et al., 2023) utilized classification results to process the self-built graph and extracted case-related information. The experimental results substantiate the beneficial impact of incorporating case-related knowledge on report generation, particularly evident when integrating real reports with the knowledge base. In addition, age and gender information are not text data and are generally encoded as one-hot vector (Zhou et al., 2021).
3.4 Multi-modal feature fusion and interaction
Feature fusion and interaction refer to the integration of multi-modal data from inputs or auxiliary tasks. This step has two purposes. First, visual and textual features from disparate domains present challenges for model learning. By fusing and facilitating interaction between these features, the domain gap can be narrowed, thereby enhancing network learning. Second, the image regions should align with the sentences in the reports. This correspondence can be learned through fusion and interaction. In the auxiliary task of embedding comparison (see Section 3.3.1), semantic features extracted from real reports are used to supervise the learning of image features. However, this approach differs from multi-modal feature fusion. The objective of embedding comparison is to enhance image features, without incorporating non-image features into the generator. Instead, the fused features in this section are forwarded to the generator.
The most straightforward approach for feature fusion and interaction is feature-level operation including the concatenation, summation, or multiplication of multimodal features (13 works). However, the feature-level operation could be too simple to enable sufficient interaction. Therefore, neural network-based methods are leveraged, such as LSTMs (2 works) and the multi-head attention mechanism (27 works). While this approach facilitates convenient feature fusion, its lack of specific design for multi-modality data fusion leads to limited effectiveness in interaction.
We would like to highlight a memory metric-based method proposed by Chen et al. (2021). It significantly enhanced the performance of the report generation system by facilitating feature interaction. A metric was initialized randomly. Image features, text features from generated tokens, and memory metric features were then projected into the same space. Subsequently, distances between image features and memory metric features, as well as text features and memory metric features, were calculated. The top K metric features with the closest distances to the image or text were selected, respectively. These selected features were then weighted based on these distances and were fed back into an encoder-decoder structure. Two studies (Qin and Song, 2022; You et al., 2022) followed this method. Wang et al. (2022a) modified it in two ways: 1) they initialized the matrix by visual and textual features; 2) the cross-modal interaction occurred only among cases with the same label. These two modifications both resulted in a notable improvement. Li et al. (2023d) contended that the methods mentioned lack explicit constraints for cross-modal alignments. They considered orthonormal bases as the metric and input them along with visual or textual features, into multi-head attention modules. Then the outputs of attention modules were processed by a self-defined gate mechanism. A triplet matching loss was utilized to align the processed visual and textual features. This method slightly improved the results.
3.5 Report generation
The last step is report generation, which utilizes extracted features from earlier steps to produce the final reports. The generation methods mainly include decoder-based techniques (Section 3.5.1), retrieval-based techniques (Section 3.5.2), and template-based techniques (Section 3.5.2). In addition, the development of large language models has made it possible to utilize them to enhance the quality of generated reports. It is discussed in Section 3.5.3
3.5.1 Decoder-based techniques
The decoder decodes the extracted representation of inputs and generates a descriptive report. The mainstream architectures include LSTM (Hochreiter and Schmidhuber, 1997) and Transformer. Compared to LSTM, the Transformer processes the entire sequence simultaneously rather than sequentially. Therefore, the Transformer allows for more efficient parallelization during training and can capture long-range dependencies. In the 89 reviewed papers, the Transformer tends to replace LSTM. Fifty-five works utilized the Transformer as a decoder and only 23 of them utilized LSTMs (12 works) or hierarchical LSTMs (11 works). For the papers published in 2023 and 2024, all encoder-decoder structures used the Transformer as their decoders. There are two ways to modify the decoder and improve model performance.
Shortcut connections: Connecting different layers in networks can be considered as a promising way to enhance the flow of information in both forward and backward propagation (Mirikharaji et al., 2023). The U-connection (Huang et al., 2023) and meshed connection (Miura et al., 2021; Lee et al., 2022; Cornia et al., 2020) are added between encoder and decoder, resulting in a similar performance enhancement (Huang et al., 2023).
Memory-driven Transformer: We would like to highlight the Memory-driven Transformer (R2Gen) proposed by Chen et al. (2020). It has been increasingly popular in recent years. The R2Gen introduces a memory module and a memory-driven conditional layer normalization module into the Transformer decoder architecture. The design of the memory module hypothesizes that diverse images exhibit similar patterns in their radiological reports, thereby serving as valuable references for each other. Building a memory matrix can capture this pattern and transfer it during training. Specifically, similar to that in the section 3.3.1, a matrix is randomly initialized and is updated using the gate mechanism based on the matrix from the last step and generated reports. The layer normalization is designed to integrate the outputs of the memory module into the decoder.
Eight works directly utilized the R2Gen as their decoder. In addition, the design of the memory module and layer normalization inspired subsequent works (Xue et al., 2024; Jia et al., 2021; Zhang et al., 2023a). It is noted that the novel utilization of the memory module by Zhang et al. (2023a) integrates ground truth reports into the training process, leading to successful outcomes.
3.5.2 Retrieval-based and template-based techniques
Retrieval-based techniques generate reports by selecting existing sentences from a large corpus and the selection is typically based on similarity comparison (Endo et al., 2021; Ramesh et al., 2022; Jeong et al., 2024). Initially, text and image encoders are trained using a contrastive method, such as the CLIP (Radford et al., 2021). The textual features of sentences in a corpus and the visual features of an input image are extracted by the encoders. The visual features are then compared with all textual features in the corpus. The top k sentences with the maximum similarity score are selected for the predicted report. In addition, Jeong et al. (2024) added a multimodal encoder after the retrieval process to calculate the image-text matching scores between the input image and the retrieved sentences. A filter was applied based on the score to remove entailed or contradicted sentences.
Other retrieval-based techniques do not follow the above process. Kong et al. (2022) treated the report generation in two steps sentence retrieval and selection. They first retrieved a candidate sentence set from the training datasets with far more sentences than a standard medical report and then selected the sentences by a classifier. Zhang et al. (2022) proposed a retrieval method based on a hashing technique, which mapped multi-modal data with the same label into a shared space.
Template-based methods typically start with the diagnosis of diseases, and then pre-defined sentences are selected based on the diagnosis results. These selected sentences are concatenated to produce reports (Pino et al., 2021). Abela et al. (2022) argued that this method was limited by exact labels, therefore they retrieved template sentences by class probabilities and different thresholds corresponding to different descriptions.
3.5.3 Large language model to assist report generation
With the emergence of ChatGPT (OpenAI, 2023), its powerful language abilities made researchers eager to harness the power of large language models to aid in report generation. However, employing it directly within the medical domain led to unsatisfactory outcomes (Tu et al., 2024; Yan et al., 2023; Sun et al., 2023). Two works (Selivanov et al., 2023; Wang et al., 2023a) initially predicted report-related information such as diseases, lesion regions, and visual features, and generated preliminary reports. Subsequently, they employed pre-trained large language models, e.g., ChatGPT (Wang et al., 2023a) or GPT-3 (Brown et al., 2020; Selivanov et al., 2023; Wang et al., 2023a) to improve the preliminary report with the predicted information and produced the final reports. The results have been moderately improved. Exploration of large language models in the field of medical report generation still requires further investigation, which is discussed in Section 7.4.
3.6 Training strategy
Training strategy refers to the techniques used to train neural network models. Traditionally, models are trained by minimizing various loss functions. Therefore, this section begins by introducing different loss functions (Section 3.6.1), followed by a discussion on reinforcement learning (Section 3.6.2), and the curriculum learning’s application to the report generation task (Section 3.6.3).
3.6.1 Loss functions
The mainstream loss function for report generation is the cross-entropy loss based on the generated sentences and the ground-truth sentences. Cross-entropy loss can be re-weighted based on term frequency (Gajbhiye et al., 2022), TF-IDF (Wang et al., 2022d) or uncertainty (Wang et al., 2024b) to mitigate model bias or handle challenging cases. In addition, Pandey et al. (2021) utilized cycle-consistency loss (Zhu et al., 2017) to generate reports. The core idea is that a report and its corresponding image share the same information, hence they can be used to generate each other.
The application of an auxiliary loss function can provide additional supervision signals, further enhancing model performance. Two works Wang et al. (2021); Li et al. (2022a) applied an additional constraint between features extracted from the generated and real reports. Zhang et al. (2023b) created two different versions of an input image by adding noise and feeding them into two networks. An auxiliary loss function ensures the consistency of the outputs produced by the two generators. Wang et al. (2024a) obtained two discriminative regions in an image from the generated words and visual classifier separately, then enforced the consistency between them.
3.6.2 Reinforcement learning
Reinforcement learning involves training an agent to make optimal decisions through trial and error, aiming to maximize targeting rewards. It offers a method to update the model parameters based on non-differentiable reward functions (Messina et al., 2022). Evaluation metrics can be considered as rewards, such as CIDEr (Kaur and Mittal, 2022b), BLEU (Qin and Song, 2022; Gu et al., 2024), METEOR (Qin and Song, 2022), ROUGE (Qin and Song, 2022), BERTScore (Miura et al., 2021), F1 score (Miura et al., 2021), and accuracy (Hou et al., 2021b). In addition, Hou et al. (2021b) trained a language fluency discriminator using the ground truth and generated reports, and then utilized the discriminator to provide rewards.
3.6.3 Curriculum learning
Liu et al. (2021a) utilized curriculum learning (Platanios et al., 2019) to classify training instances and trained a model from simple to complex samples. Data pairs were evaluated based on image heuristics, image confidence, text heuristics, and text confidence. Image heuristic evaluated the similarity between input images and normal images. Image confidence indicated the confidence of a classification model. The report heuristic was related to the number of abnormal sentences, and report confidence was evaluated by the negative log-likelihood loss (Xu et al., 2020).
4 Datasets
Datasets play a crucial role in the development of report-generation models. Abundant and diverse training data can improve the model’s accuracy and generalizability. Moreover, a suitable test dataset makes it realistic to test the model’s performance in a practical scenario. In this section, we selected 11 public medical image-report datasets utilized in the reviewed articles to provide a comprehensive introduction of popular and newly collected datasets, see Table 1. The usage of datasets in each article is shown in Table LABEL:tab:overall in Appendix A. The datasets primarily focus on the lung, including X-ray (Demner-Fushman et al., 2016; Johnson et al., 2019a, b) and CT (Li et al., 2020; Liu et al., 2021e). In addition, there are also publicly available datasets on eye scans (Huang et al., 2021b; Lin et al., 2023; Li et al., 2021) and breast scans (Yang et al., 2021a). The concentration of medical report generation efforts on chest X-rays can be attributed, in part, to the accessibility of large-scale publicly available datasets.
Figure 4: A sample from the MIMIC-CXR dataset. The left is the Chest X-ray image and the right is its corresponding radiology report.
The most popular datasets are the Indiana University Chest X-Ray Collection (IU X-Ray) (Demner-Fushman et al., 2016) and the MIMIC Chest X-ray (MIMIC-CXR) (Johnson et al., 2019a, b). The IU X-ray contains 7470 images of frontal and lateral X-rays and 3955 reports with manual annotation based on the MeSH codes [^3] https://www.nlm.nih.gov/mesh/meshhome.html and the RadLex codes. These codes encompass standard medical terminology, such as anatomical structures, diseases, pathological signs, foreign objects, and attributes, as defined by authoritative institutions. For deep learning methods, the size of the IU X-RAY is insufficient, while the collection of MIMIC-CXR alleviates this problem (Messina et al., 2022). It consists of 377,110 images and 227,827 reports, and has been multi-labeled by automatic tools according to 14 disease categories. An image-report sample from the MIMIC-CXR dataset is shown in the Figure 4. The indication, technique, and comparison provide fundamental information for a Chest X-ray test. The automation of radiology report generation targets the findings and impressions sections. The findings section provides a detailed description of the entire image, while the impressions section summarizes these observations. The follow-up work provides MIMIC-CXR with richer label information. The Chest ImaGenome dataset (Wu et al., 2021) is based on the anteroposterior and posteroanterior view Chest X-ray images in the MIMIC-CXR dataset. It provides an anatomy-centered scene graph for each image, including anatomical location and relation annotations. The annotations are generated through two automated pipelines, and a separate dataset of 500 manually annotated cases is also provided for testing.
Increasing attention has been paid to the collection of ophthalmic image-report pairs (Huang et al., 2021b; Li et al., 2021; Lin et al., 2023). Ophthalmic image has lots of modalities, such as Fluorescein Angiography (FA), Fundus Fluorescein Angiography (FFA), Color Fundus photography (CFP), fundus photograph (FP), optical coherence tomography (OCT), fundus autofluorescence (FAF), Indocyanine Green Chorioangiography (ICG), and red-free filtered fundus images. Most of retinal datasets are based on one or two modalities and the last released Retina ImBank dataset is the first multi-modality retinal image-text dataset (Lin et al., 2023). Some datasets provide additional labels for each image, such as lesion boundary (Li et al., 2021), lesion category (Huang et al., 2021b; Li et al., 2021), and keywords (Huang et al., 2021b).
All the above datasets provide unstructured reports. Recently, Pellegrini et al. (2023) released a structured report dataset named Rad-ReStruct based on the IU-XRAY. In clinical practice, generating structured reports typically requires doctors to answer a sequence of questions (Pellegrini et al., 2023). Therefore, Pellegrini et al. (2023) designed a structured report template with a series of single- or multi-choice questions based on topic existence (e.g., Are there any diseases in the lung?), element existence (e.g., Is there an opacity in the lung?), and attributes (e.g., What is the degree?). They then integrated the IU-XRAY report data into the template using its annotated MeSH and RadLex codes.
Table 1: Public datasets for medical report generation.
Name Image type Report type Images Reports Patients Used by
IU X-Ray (Demner-Fushman et al., 2016) Chest X-ray Unstructured 7,470 3,955 3,955 67 works
MIMIC-CXR (Johnson et al., 2019a, b) Chest X-ray Unstructured 377,110 227,827 65,379 59 works
Chest ImaGenome (Wu et al., 2021) Chest X-ray Unstructured 242,072 217,013 4 works
COV-CTR (Li et al., 2020) Chest CT Unstructured 728 728 4 works
DeepEyeNet (Huang et al., 2021b) FA, CFP Unstructured 15,709 15,709 3 works
FFA-IR (Li et al., 2021) FFA Unstructured 1,048,584 10,790 2 works
Chinese COVID-19 CT (Liu et al., 2021e) Chest CT Unstructured 1,104 368 96 1 work
BCD2018 (Yang et al., 2021a) Breast Ultrasound Unstructured 5,349 5,349 1 work
Retina ImBank (Lin et al., 2023) FP, OCT, FFA, FAF, ICG, and red-free filtered fundus images Unstructured 18,788 18,788 1 work
Retina Chinese (Lin et al., 2023) FP, FFA, and ICG Unstructured 57,498 57,498 1 work
Rad-ReStruct (Pellegrini et al., 2023) Chest X-ray Structured 3,720 3,597 3,597 1 work
5 Evaluation
Accurate assessment of the quality of generated reports is crucial for measuring model performance. The quality of generated reports can be evaluated both quantitatively and qualitatively. Quantitative methods check the text quality and medical correctness of the generated report by natural language evaluation metrics (Section 5.1) and medical correctness metrics (Section 5.2), respectively. Qualitative evaluation is normally performed by human experts and they provide overall evaluation for the generated reports (Section 5.3). Table 2 summarizes the description of evaluation methods introduced in this section. The usage of evaluation methods in each article is shown in Table LABEL:tab:overall in Appendix A.
Table 2: The quantitative and qualitative evaluation methods for medical report generation.
Metrics Description Used by
Natural language-based metrics
BlEU
(Papineni et al., 2002)
A precision-based metric that measures the n-gram overlapping of the generated text and ground truth text. 84 works
ROUGE-L
(Lin, 2004)
A F1-like metric that computes a weighted harmonic mean of precision and recall based on the longest common subsequence. 75 works
METEOR
(Banerjee and Lavie, 2005)
A F1-like metric that computes a weighted harmonic mean of unigram precision and recall. It is an extension of BLEU-1. 57 works
CIDEr
(Vedantam et al., 2015)
The cosine similarity between generated text and ground truth text based on the TF-IDF. 41 works
$S_{emb}$
(Endo et al., 2021)
Sending ground truth report and generated report into a textual feature extractor and calculating the cosine similarity between their embeddings from the last layer. 2 works
%Novel
(Van Miltenburg et al., 2018)
The percentage of generated descriptions that are not present in the training data. 1 work
Medical correctness-based metrics
Clinical Efficacy
(Chen et al., 2020; Liu et al., 2019)
Calculate accuracy, precision, recall, and F1 score based on observations extracted from reference reports and generated reports by automated system. 34 works
MIRQI
(Zhang et al., 2020)
Calculate precision, recall, and F1 score based on graph comparison. The ground truth and generated reports are automatically analysed to construct a sub-graph from a defined abnormality graph. 2 work
nKTD
(Zhou et al., 2021)
Calculate the Hamming distance based on observations extracted from reference reports and generated reports by the CheXpert Labeller. 1 work
Human-based evaluation
Comparison Generate reports by two different models and allow senior radiologists to find which report is better. 12 works
Classification Radiologists categorize the produced reports as accurate, missing details, and false reports. 1 work
Error scoring Radiologists assess the error severity of baseline, model generated, and reference reports. 1 work
Grading Radiologists need to assign a 5-point scale grade to two types of generated reports, in accordance with clinical standards. 1 work
5.1 Natural language-based evaluation metrics
Natural language evaluation metrics are from natural language processing tasks and measure the general text quality of generated reports. In the reviewed papers, the most popular metrics are BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and CIDEr (Vedantam et al., 2015), which are based on n-gram matching between reference reports and generated reports. The model is deemed superior with an increased number of matches. Among them, the BLEU is the earliest and proposed a modified precision method. When evaluating the quality of radiology report generation, we typically opt for BLEU-1, BLEU-2, BLEU-3, and BLEU-4 metrics. The n in BLEU-n means the calculation is based on n-gram. The METEOR is an extension of BLEU-1 and introduces recall into evaluation. The ROUGE-L also considers precision and recall based on the longest common subsequence between reference and generated text. The CIDEr adopts the TF-IDF. The TF-IDF vectors weigh each n-gram in a sentence, and then the cosine similarity is calculated between the TF-IDF vectors of reference and generated text. When the model consistently produces the most common sentences, it can achieve notable BLEU scores. However, CIDEr can evaluate generated outputs by encouraging the appearance of important terms and punishing high-frequency vocabulary (Li et al., 2023a). CIDEr has a popular variant CIDEr-D, which introduces penalties to generate desired sentence length and remove stemming to ensure the proper usage of word forms.
Other than n-gram matching, Endo et al. (2021) proposed a new metric $S_{emb}$. A pre-trained feature extractor is applied on both the ground truth and the generated reports, and the cosine similarity between extracted embeddings is calculated. This approach is used to assess whether the semantic information contained in two sentences is consistent. Another natural language metric %Novel (Van Miltenburg et al., 2018) is introduced to evaluate the diversity in image captioning.
5.2 Medical correctness metrics
Natural language evaluation metrics evaluate the similarity between produced reports and the ground truth, but cannot accurately measure whether the generated reports contain the required medical facts (Babar et al., 2021b; Messina et al., 2022). So medical correctness metrics are proposed to pay attention to the prediction of important medical facts. Generally, an automatic labeller is applied to extract medical facts from generated and reference reports. Then different metrics are applied to these. The mainstream metric is clinical efficacy (Liu et al., 2019; Chen et al., 2020), which initially calculates precision, recall, and F1 score, and subsequently extends to accuracy (Babar et al., 2021a; Miura et al., 2021; Moon et al., 2022; Yan, 2022; Yang et al., 2022; Selivanov et al., 2023; Yang et al., 2023) and AUC (Li et al., 2023a). Most of the works utilized CheXpert (Irvin et al., 2019) as a labeler to extract chest diseases information from ground truth reports, while seven of them (Miura et al., 2021; Yan, 2022; Wang et al., 2023a, 2024a; Dalla Serra et al., 2023b, a; Jin et al., 2024) utilized a newer labeler CheXbert (Smit et al., 2020), which has a higher performance. In addition, Pellegrini et al. (2023) generated structured reports by predicting a series of questions. They utilized macro precision, recall, and F1 score to assess all questions, along with evaluating report-level accuracy. Other metrics calculate the Hamming distance (Zhou et al., 2021) or perform graph comparison (Zhang et al., 2020) based on the extracted results.
5.3 Human-based evaluation
For qualitative assessment, the most common human-based evaluation method is comparison. In general, a set number of samples (i.e., 100/200/300) are selected from the test dataset and subsequently processed by different generated models. More than one professional clinician is responsible to compare and sort the generated reports. Five works utilized ground truth reports in this process. Two works (Miura et al., 2021; Cao et al., 2022) considered the ground truth reports as a reference. Reports were generated by different generators and the radiologists need to select which report is more similar to the reference. Dalla Serra et al. (2022) allowed experts to find 5 types of errors (i.e., hallucination, omission, attribute error, impression error, and grammatical error) in different generated reports according to the reference reports. Xu et al. (2023) asked radiologists to rank the ground truth reports and the generated reports. Qin and Song (2022) invited experts to select the most suitable report from the generated and the ground truth reports according to correctness, language fluency, and content coverage. Other expert evaluation methods are classification (Alfarghaly et al., 2021), grading (Wang et al., 2024b), and error scoring (Jeong et al., 2024).
6 Benchmark Comparison
For model comparison, it is essential to select a benchmark for an impartial evaluation. We choose the MIMIC-CXR (Johnson et al., 2019b, a)(see Table 3) dataset as a benchmark to compare the model performance for two reasons. First, according to Table 1, IU-Xray and MIMIC-CXR are the most popular datasets, but IU X-ray lacks standard training-validation-test splits, leading to less comparable results (Messina et al., 2022), while MIMIC-CXR has official training-validation-test splits. Second, MIMIC-CXR is the largest image-report dataset, which provides a broader distribution of data, facilitating testing across diverse scenarios, and reducing biases commonly encountered in smaller datasets.
We endeavour to ensure equitable comparisons, but it’s important to consider the following three problems when analysing these results. First, we select the methods leveraged the official splits of the MIMIC-CXR dataset (Johnson et al., 2019b), containing 368960 images (with 222758 reports) in the training dataset, 2991 images (with 1808 reports) in the validate dataset, and 5159 images (with 3269 reports) in the test dataset. Although some papers claimed to use official splits, the total number of datasets is different. For example, the MIMIC-CXR used by Najdenkoska et al. (2022) contains 218,101 samples. These papers were excluded. In Section 3, a multitude of methodologies are highlighted. However, due to stringent screening, their absence from the comparison table does not imply inferior performance. Second, we choose BLEU, ROUGE, METEOR, CIDEr-D, precision, recall, and F1 score metrics as evaluation metric. Similar to the previous survey (Messina et al., 2022), we found that although some metrics have variants, many papers do not specify the particular version used. In that case, we assume they are consistent. Third, the generated report sections vary among different methodologies. Most articles do not explicitly specify the generated report sections, making it challenging to conduct comparative analysis.
We separate the comparison results based on the generated report section, and the completed comparison tables are provided in Table A2 in Appendix A. Table 3 shows the best and second best performances ranked for each evaluation metric. Among the 14 papers in Table 3, three of them (Song et al., 2022; Liu et al., 2023b; Jin et al., 2024) utilized multi-modality inputs. From an architectural perspective, in the encoding stage, five works utilized a pure Transformer model (Wang et al., 2022e; Kong et al., 2022; Liu et al., 2023b; Wang et al., 2023d, 2022d), another five works employed the CNN-based model (Pino et al., 2021; Song et al., 2022; Wu et al., 2022; Jia et al., 2022; Jin et al., 2024), and four works combined CNN and Transformer (Chen et al., 2021; Wang et al., 2023c, 2024a; Wu et al., 2023). In the generation stage, most papers relied on the Transformer. In terms of technical modules, six works (Wang et al., 2022e; Jin et al., 2024; Wang et al., 2023d; Wu et al., 2023; Wang et al., 2024a, 2022d) designed auxiliary tasks and five works (Song et al., 2022; Wu et al., 2022; Wang et al., 2023d; Liu et al., 2023b; Wang et al., 2022e) used contrastive learning.
Table 3: Comparisons of the model performance on the MIMIC-CXR Dataset. B1, B2, B3, B4, R-L, C-D, P, R, and F represent BLEU-1, BLEU-2, BLEU-3, BLEU-4, ROUGE-L, CIDEr-D, precision, recall, and F1 score, respectively. The best and second best results are highlighted. All values were extracted from their papers.
[…]
Findings Section
Chen et al. (2021) 0.353 0.218 0.148 0.106 0.278 0.142 - 0.334 0.275 0.278
Pino et al. (2021) - - - - 0.185 - 0.238 0.381 0.531 0.428
Song et al. (2022) 0.360 0.227 0.156 0.117 0.287 0.148 - 0.444 0.297 0.356
Impression + Findings Section
Wu et al. (2022) 0.340 0.212 0.145 0.103 0.270 0.139 0.109 - - -
Wang et al. (2022d) 0.351 0.223 0.157 0.118 0.287 - 0.281 - - -
Jia et al. (2022) 0.363 0.228 0.156 0.130 0.300 - - - - -
Unspecified generated sections
Wang et al. (2023c) 0.363 0.235 0.164 0.118 0.301 0.136 - - - -
Wang et al. (2024a) 0.374 0.230 0.155 0.112 0.279 0.145 0.161 0.483 0.323 0.387
Wu et al. (2023) 0.383 0.224 0.146 0.104 0.280 0.147 - - - 0.758
Wang et al. (2023d) 0.386 0.250 0.169 0.124 0.291 0.152 0.362 0.364 0.309 0.311
Liu et al. (2023b) 0.391 0.249 0.172 0.125 0.304 0.160 - - - -
Jin et al. (2024) 0.398 - - 0.112 0.268 0.157 - 0.501 0.509 0.476
Wang et al. (2022e) 0.413 0.266 0.186 0.136 0.298 0.170 0.429 - - -
Kong et al. (2022) 0.423 0.261 0.171 0.116 0.286 0.168 - 0.482 0.563 0.519
7 Challenges and Future Works
Although automated systems offer promising efficiency for clinical workflows, current methods have not produced very high-quality reports. This section evaluates the current progress in automated report generation development and identifies potential areas for improvement.
7.1 Constructing and utilizing multi-modal data
Considering report generation as a multi-modal problem is more aligned with clinical practice (Tu et al., 2024; Yan et al., 2023). Babar et al. (2021a) have proven the ineffectiveness of simple encoder-decoder report generation models and mentioned that adding prior knowledge can be a promising method. However, the current utilization of multi-modal data remains under-explored. Firstly, the methods for non-image feature extraction and the fusion of multi-modality data are often limited and simplistic, such as using graph encoding for the knowledge base and attention mechanisms for feature fusion. Secondly, the construction of the knowledge base is imperfect. The pre-defined graph (Zhang et al., 2020) is overly simplistic. Despite Radgraph being a vast knowledge base, it is solely derived from reports, lacking the relationship between images and reports, such as organ recognition or understanding of typical radiological scenarios, which radiologists possess. The Chest ImaGenome dataset (Wu et al., 2021) provides the organ recognition annotations, alleviating the problem. Additionally, as far as we know, publicly available knowledge bases only concentrate on chest X-rays, leaving a gap in general medical knowledge databases.
7.2 Evaluation of medical correctness
Evaluating the medical correctness of generated reports is crucial for the clinical application. Compared to previous works (Messina et al., 2022; Liao et al., 2023), recent works have paid more attention to it, but still have two shortcomings. First, in the reviewed articles, medical correctness evaluation has only been applied to chest X-ray reports. Secondly, the evaluation is based on the automatic labeler of radiology reports, which are only targeted at 14 types of diseases and the average F1 score is around 0.798 (Smit et al., 2020). Thus, improving the accuracy and scale of automatic labeling tools can help optimize the evaluation process.
7.3 Large public datasets and unified comparison benchmark
As shown in Table 1, most of the public datasets are limited in size. Deep learning-based techniques require a large amount of data. The contemporary prevalence of large language models underscores this need for extensive data volumes. Among the datasets, the MIMIC dataset is relatively large but only includes Chest X-ray data. Large datasets targeting other image modalities and diseases need to be constructed. In addition, while MIMIC-CXR is a well-established benchmark compared to IU-XRay, dataset utilization lacks standardization, complicating comparisons. We urge papers using the MIMIC dataset to define their training, validation, and testing partitions, with explicit disclosure of the filtration method, particularly for the testing dataset.
7.4 Human-AI interaction
Most papers overlook the interaction between users (e.g., clinicians or patients) and automated systems. When the system performs as an AI assistant, users may want to know the insights of the model regarding specific aspects of medical images. In the reviewed works, Tanwani et al. (2022) constructed a Visual Question Answering system for medical report generation, and Tanida et al. (2023) linked the output results to image regions using object detection, allowing users to select areas of interest and receive corresponding language explanations. Recently, dialogue systems (e.g., GPT-4 (Achiam et al., 2023), PaLM (Chowdhery et al., 2023), Gemini (Team et al., 2023)) based on large language models and large multi-modal models have shown people more possibilities for human-AI interaction. Although general large language models can provide answers to some questions in medical question answering benchmarks, their deployment in clinical settings remains unfeasible due to safety concerns within the medical domain (Yan et al., 2023; Saab et al., 2024).
To enable dialogue systems to comprehend medical knowledge, fine-tuning the model with medical data is an intuitive approach, such as LLaVA-Med (Li et al., 2024) and Med-PaLM 2 (Singhal et al., 2023). However, they did not test the model’s performance on the report generation task. Saab et al. (2024) proposed Med-Gemini, a series of highly proficient multi-modal models tailored specifically for the medical domain. They intuitively showed the interactive report generation process for a normal case, while lacking quantitative results to evaluate report generation performance. Exploring the direction of report generation with human-AI interaction holds significant promise. Additionally, fine-tuning large models demands substantial GPU resources, making efficient methods crucial.
7.5 Standardized report generation
Most of the existent methods focused on unstructured report generation, while structured reporting has several advantages, such as saving time (Hong and Kahn, 2013; Nobel et al., 2022), preventing errors, decreasing communication expenses linked to ambiguous natural language. Recently, Pellegrini et al. (2023) developed a structured template and released a related dataset. This promising start could pave the way for further exploration in this direction.
8 Conclusions
In this study, we have conducted a detailed technical review of 89 papers on automatic medical report generation published in the years 2021, 2022, 2023, and 2024 to showcase both mainstream and novel techniques. Our particular focus lies on the utilization and fusion of multi-modality data. The analysis of methods is structured based on the components of the report generation pipeline, presenting the key techniques for each component. Additionally, we provide a summary of the current publicly available datasets and evaluation methods, encompassing both quantitative and qualitative assessments. Subsequently, we compare the results from a subset of the 89 papers under the same experimental setting. Finally, we outline the current challenges and propose future directions for medical report generation. Overall, sustained progress is needed to produce standardized and clinically accurate reports. This survey aims to offer a comprehensive overview of report-generation techniques, emphasize critical issues, and assist researchers in promptly grasping recent advancements in the field to build more robust systems for clinical practice.
Appendix A
Table A1: Summary of papers in the survey. The overview is based on the findings of the survey. The dataset and metrics section focuses on mainstream datasets and metrics. ’–’ indicates that the paper either does not detail this process or does not include the key techniques summarized in this survey. The following abbreviations are used: I-Archi: the architecture of image feature learning, I-Module: the enhancement module of image feature learning, NI: the feature learning of non-image data, GEI: gastrointestinal endoscope image, RetiI: retinal image, Term: terminology, KnowB: knowledge base, RealR: real report, ClinicalI: clinical information, Ques: questionnaires, FreFilter: frequency-based filtering, ConAtoL: converting all tokens to lowercase, RemoNAT: removing non-alphabetic tokens, AT: auxiliary task, ContrasL: Contrastive learning, MM: memory metric, FeatO: feature-level operation, OptimS: optimization strategies, H-LSTM: hierarchical LSTM, ReLoss: re-weighted loss function, ReinL: reinforcement learning, R-L: Rouge-L, C-D: CIDEr-D, CE: clinical efficacy, Com: comparison, Clas: classification, Escore: error scoring, ECI: extracting case-related information, IU: IU X-Ray, MIMIC: MIMIC-CXR, ImaGeno: Chest ImaGenome, COV: COV-CTR, DeepEye: DeepEyeNet, CCCT: Chinese COVID-19 CT, Ret-I: Retina ImBank, and Ret-C: Retina Chinese. In addition, AT-Graph, AT-Class, AT-EC, and AT-DS mean the graph-based, classification, embedding comparison, and detection/segmentation auxiliary tasks.
Paper Input data Data preparation Feature Learning Feature Fusion Generation Training Strategy Datasets Metrics
I-Archi I-Module NI
Liu et al. [2021b] Chest X-ray,
Term,
RealR
Tokenizing,
ConAtoL,
RemoNAT,
FreFilter
ResNet AT-Graph Transformer FeatO Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D
Liu et al. [2021d] Chest X-ray ResNet AT-Graph Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
CE,
Com
Liu et al. [2021a] Chest X-ray Tokenizing,
ConAtoL,
RemoNAT
CNNs LSTMs Curricu-
lum lea-
rning
IU,
MIMIC
BLEU,
R-L,
METEOR,
Com
Chen et al. [2021] Chest X-ray ResNet+
Transformer
MM Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
CE
You et al. [2021] Chest X-ray ResNet Attention Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
Com
Miura et al. [2021] Chest X-ray DenseNet+
Transformer
Transformer ReinL MIMIC BLEU,
C-D,
CE,
Com
Alfarghaly et al. [2021] Chest X-ray Resizing DenseNet AT-Class Attention Transformer IU BLEU,
R-L,
METEOR,
C-D,
Clas
Yang et al. [2021b] Chest X-ray ResNet AT-EC H-LSTM IU,
MIMIC
BLEU,
R-L,
C-D
Pahwa et al. [2021] Chest X-ray Resizing,
Cropping,
Flipping
HRNet+
Transformer
R2Gen IU BLEU,
R-L,
METEOR
Zhou et al. [2021] Chest X-ray,
ClinicalI
Tokenizing,
FreFilter
DenseNet AT-Class,
AT-EC
One-hot,
BioSentVec
Attention H-LSTM IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D,
nKTD
Huang et al. [2021b] RetiI,
Term
CNNs Embedding
layer
FeatO LSTM DeepEye BLEU,
R-L,
C-D
Han et al. [2021] Spine MRI Self-design AT-DS Reasoning
Liu et al. [2021e] Chest X-ray,
Chest CT,
Term
DenseNet AT-Class BERT FeatO,
Attention
Transformer AT CCCT BLEU,
R-L,
C-D,
Com
Wang et al. [2021] Chest X-ray,
Chest CT
Tokenizing ResNet+
Transformer
AT-EC H-LSTM AT IU,
COV
BLEU,
R-L,
C-D
Endo et al. [2021] Chest X-ray ResNet Retrieval MIMIC BLEU,
$S_{emb}$,
CE
Najdenkoska et al. [2021] Chest X-ray Resizing,
Tokenizing,
ConAtoL,
RemoNAT,
FreFilter
DenseNet+
Transformer
AT-EC LSTM IU,
MIMIC
BLEU,
R-L,
METEOR,
CE
Yang et al. [2021a] Breast
ultrasound
Tokenizing,
FreFilter
ResNet AT-Class FeatO LSTM BCD2018 BLEU,
R-L,
METEOR,
C-D
Pandey et al. [2021] Chest X-ray Resizing VGG H-LSTM AT IU BLEU,
R-L
Hou et al. [2021b] Chest X-ray Resizing,
Tokenizing,
FreFilter
ResNet AT-Class LSTM H-LSTM ReinL IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D,
CE
Pino et al. [2021] Chest X-ray DenseNet Template IU,
MIMIC
BLEU,
R-L,
C-D,
CE,
MIRQI
Liu et al. [2021c] Chest X-ray Tokenizing,
ConAtoL,
FreFilter
ResNet ContrasL H-LSTM IU,
MIMIC
BLEU,
R-L,
METEOR,
CE,
Com
Jia et al. [2021] Chest X-ray ResNet+
Transformer
Transformer IU,
MIMIC
BLEU,
R-L
Hou et al. [2021a] Chest X-ray Resizing,
Data
augmentation
DenseNet Transformer MIMIC BLEU,
R-L,
METEOR,
CE
Huang et al. [2021a] RetiI,
Term
Tokenizing,
ConAtoL,
RemoNAT,
FreFilter
CNNs Embedding
layer
LSTM LSTM DeepEye BLEU,
R-L,
C-D
Babar et al. [2021a] Chest X-ray Tokenizing,
ConAtoL,
RemoNAT
Uncondition IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D,
CE
Singh et al. [2021] Chest X-ray Resizing,
Tokenizing,
ConAtoL,
RemoNAT,
FreFilter
InceptionV3 AT-Class LSTM IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D
Yang et al. [2022] Chest X-ray,
KnowB,
RealR
Resizing,
Tokenizing,
ConAtoL,
FreFilter
ResNet RotatE,
ECI,
BERT
Attention Transformer IU,
MIMIC
BLEU,
R-L,
C-D,
CE
Ramesh et al. [2022] Chest X-ray Tokenizing, Filtering ResNet/
Transformer
Retrieval MIMIC $S_{emb}$
Sirshar et al. [2022] Chest X-ray Tokenizing,
RemoNAT,
ConAtoL
VGG LSTM IU,
MIMIC
BLEU
Najdenkoska et al. [2022] Chest X-ray Resizing,
RemoNAT,
FreFilter
DenseNet+
Transformer
AT-EC Transformer,
LSTM
IU,
MIMIC
BLEU,
R-L,
METEOR,
%Novel,
CE
Wang et al. [2022c] Chest X-ray Cropping,
Tokenizing,
ConAtoL,
FreFilter
DenseNet AT-Graph H-LSTM IU BLEU,
R-L,
C-D
Li et al. [2022b] RetiI Resizing,
Tokenizing,
ConAtoL,
FreFilter
I3D+
Transformer
AT-Graph Transformer FFA-IR BLEU,
R-L,
METEOR,
C-D,
Com
Qin and Song [2022] Chest X-ray ResNet+
Transformer
MM Transformer ReinL IU,
MIMIC
BLEU,
R-L,
METEOR,
CE,
Com
Wang et al. [2022a] Chest X-ray Resizing,
Cropping
ResNet+
Transformer
ContrasL MM Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D
Cao et al. [2022] Chest X-ray,
GEI,
Term
DenseNet+
Transformer
AT-Graph BERT FeatO,
Attention
Transformer IU BLEU,
R-L,
C-D,
Com
Mohsan et al. [2022] Chest X-ray Tokenizing,
ConAtoL
Transformer Transformer IU BLEU,
R-L,
METEOR,
C-D
Yan [2022] Chest X-ray DenseNet+
Transformer
AT-Graph,
MM
Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D,
CE,
MIRQI
Li et al. [2022a] Chest X-ray ResNet+
Transformer
AT-Class Transformer AT IU BLEU,
R-L,
METEOR
Sun et al. [2022] RetiI Resizing ResNet+
Faster-RCNN
AT-EC,
AT-DS
Transformer FFA-IR BLEU,
R-L,
METEOR,
C-D
You et al. [2022] Chest X-ray Tokenizing,
ConAtoL,
RemoNAT
ResNet+
Transformer
AT-Class Attention,
MM
Transformer IU BLEU,
R-L,
METEOR
Wang et al. [2022d] Chest X-ray Resizing Transformer AT-Class,
AT-EC
Transformer ReLoss IU,
MIMIC
BLEU,
R-L,
C-D
Lee et al. [2022] Chest X-ray Resizing,
Cropping,
Flipping
ResNet+
Transformer
R2Gen IU BLEU,
R-L,
METEOR
Moon et al. [2022] Chest X-ray Resizing,
Cropping,
Tokenizing
ResNet+
Transformer
AT Transformer IU,
MIMIC
BLEU,
CE
Huang et al. [2022] RetiI,
Term
Resizing,
Tokenizing,
ConAtoL,
RemoNAT,
FreFilter
CNNs Embedding
layer
Attention LSTM DeepEye BLEU,
R-L,
C-D
Kaur and Mittal [2022a] Chest X-ray Resizing,
Convert image to
grayscale,
Flipping,
Tokenizing,
ConAtoL,
RemoNAT
VGG H-LSTM IU BLEU,
R-L,
C-D
Song et al. [2022] Chest X-ray,
RealR
DenseNet ContrasL Attention Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
CE
Jia et al. [2022] Chest X-ray DenseNet Attention H-LSTM IU,
MIMIC
BLEU,
R-L
Zhang et al. [2022] Chest X-ray Resizing VGG Retrieval MIMIC Precision
Dalla Serra et al. [2022] Chest X-ray,
ClinicalI
Tokenizing,
Resizing,
Flipping,
Rotation,
Cropping
ResNet+
Transformer
AT-Graph Embedding
layer
Attention Transformer MIMIC BLEU,
R-L,
METEOR,
CE,
Com
Yan et al. [2022] Chest X-ray Tokenizing,
ConAtoL,
FreFilter
ResNet+
Transformer
AT Attention Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D,
CE
Gajbhiye et al. [2022] Chest X-ray Tokenizing,
Removing irrelevant elements, ConAtoL,
FreFilter
DenseNet AT-Class LSTM ReLoss IU BLEU,
R-L,
METEOR,
C-D
Wang et al. [2022b] Chest X-ray Grouping ResNet+
Transformer
AT-Class R2Gen IU,
MIMIC
BLEU,
R-L,
METEOR
Abela et al. [2022] Chest X-ray DenseNet Template MIMIC BLEU,
CE
Wu et al. [2022] Chest X-ray ResNet ContrasL LSTM IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D
Tanwani et al. [2022] Chest X-ray,
Ques
Resizing,
Image transformations,
Tokenizing
ResNeXt AT-Class,
ContrasL
BERT Attention Transformer IU BLEU
Chen et al. [2022] Chest X-ray Resizing,
Tokenizing,
ConAtoL,
RemoNAT
ResNet+
Transformer
AT-EC
MM
R2Gen IU BLEU,
R-L,
METEOR
Kong et al. [2022] Chest X-ray Resizing Transformer Retrieval IU,
MIMIC
BLEU,
R-L,
METEOR,
CE
Wang et al. [2022e] Chest X-ray Tokenizing,
ConAtoL,
RemoNAT,
FreFilter
Transformer AT-Class,
ContrasL
FeatO Transformer MIMIC BLEU,
R-L,
METEOR,
C-D
Du et al. [2022] Chest X-ray ResNet AT-Class H-LSTM IU BLEU,
R-L,
METEOR
Kaur and Mittal [2022b] Chest X-ray Resizing,
Tokenizing,
ConAtoL,
RemoNAT,
FreFilter
VGG AT-Class FeatO H-LSTM ReinL IU BLEU,
R-L,
C-D
Tanida et al. [2023] Chest X-ray Resizing,
Data
augmentation,
Removing redundant whitespaces
ResNet+
Faster-RCNN
AT-DS Transformer MIMIC,
ImaGeno
BLEU,
R-L,
METEOR,
C-D,
CE
Wang et al. [2023d] Chest X-ray Tokenizing Transformer AT Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D,
CE
Li et al. [2023b] Chest X-ray,
Chest CT
Resizing,
Tokenizing,
FreFilter
DenseNet AT-Graph Attention Transformer IU,
COV
BLEU,
R-L,
C-D,
Com
Wu et al. [2023] Chest X-ray Tokenizing,
ConAtoL,
Removing irrelevant elements,
FreFilter
ResNet+
Transformer
Transformer ReinL,
AT
IU,
MIMIC
BLEU,
R-L,
METEOR
Li et al. [2023a] Chest X-ray,
KnowB,
RealR
Tokenizing Transformer AT-Class,
ContrasL
ECI,
BERT
Attention Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D,
CE
Huang et al. [2023] Chest X-ray,
KnowB
Tokenizing,
ConAtoL,
RemoNAT,
FreFilter
ResNet+
Transformer
ECI,
BERT
FeatO,
Attention
Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
CE
Zhang et al. [2023b] Chest X-ray DenseNet+
Transformer
AT-Graph,
AT-Class,
AT-EC
Transformer AT IU,
MIMIC
BLEU,
R-L,
METEOR,
CE
Xu et al. [2023] Chest X-ray,
Dermoscopy,
KnowB
DenseNet+
Transformer
AT-Class ECI,
BERT
Attention R2Gen IU BLEU,
R-L,
METEOR,
Com
Yang et al. [2023] Chest X-ray,
KnowB
Resizing,
Tokenizing,
ConAtoL,
FreFilter
ResNet AT-Class,
AT-EC
Attention Transformer IU,
MIMIC
BLEU,
C-D,
CE
Zhang et al. [2023a] Chest X-ray,
Chest CT
Resizing ResNet+
Transformer
Transformer+
MM
IU,
MIMIC,
COV
BLEU,
R-L,
METEOR
Wang et al. [2023a] Chest X-ray ResNet+
Transformer
AT-Class,
AT-DS
R2Gen
+LLM
MIMIC CE
Lin et al. [2023] Chest X-ray,
RetiI
Resizing,
Tokenizing,
FreFilter
ResNet+
Transformer
ContrasL Transformer IU,
MIMIC,
Ret-I,
Ret-C
BLEU,
R-L,
METEOR
Selivanov et al. [2023] Chest X-ray Resizing,
Tokenizing,
ConAtoL,
RemoNAT
DenseNet Transformer IU,
MIMIC
BLEU,
R-L,
C-D,
CE
Cao et al. [2023] Chest X-ray,
GEI,
Term
DenseNet+
Transformer
BERT FeatO,
Attention,
MM
Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D
Shetty et al. [2023] Chest X-ray Self-design LSTM IU BLEU
Wang et al. [2023c] Chest X-ray - ResNet+
Faster-RCNN+
Transformer
R2Gen IU,
MIMIC,
ImaGeno
BLEU,
R-L,
METEOR
Li et al. [2023d] Chest X-ray Resizing,
Cropping,
Tokenizing,
FreFilter,
ConAtoL,
RemoNAT
Transformer MM
+Self-design
Transformer AT IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D
Liu et al. [2023b] Chest X-ray,
Term,
RealR
FreFilter,
ConAtol
DenseNet ContrasL Transformer FeatO,
Attention
Transformer IU,
MIMIC
BLEU,
R-L,
METEOR
Wang et al. [2023b] Chest X-ray ResNet+
Transformer
ContrasL R2Gen IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D
Pellegrini et al. [2023] Chest X-ray,
Ques
EfficientNet BERT Attention Classification - Rad-ReStruct CE
Dalla Serra et al. [2023a] Chest X-ray,
ClinicalI
Resizing,
Cropping,
Grouping
ResNet+
Faster-RCNN
AT-DS Embedding
layer
Attention Transformer MIMIC,
ImaGeno
BLEU,
R-L,
METEOR,
CE
Li et al. [2023c] Chest X-ray,
RealR,
Term
Resizing,
Self-design
ResNet+
Transformer
- ECI,
BERT
Attention Transformer - IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D
Dalla Serra et al. [2023b] Chest X-ray,
ClinicalI
Resizing,
Cropping
ResNet+
Faster-RCNN
AT-DS,
AT-Graph
Embedding
layer
Attention Transformer - MIMIC,
ImaGeno
BLEU,
R-L,
METEOR,
CE
Jeong et al. [2024] Chest X-ray Resizing Transformer Retrieval MIMIC BLEU,
Escore
Gu et al. [2024] Chest X-ray,
Term
Resizing,
Cropping,
Flipping,
FreFilter,
RemoNAT
ResNet AT-DS,
AT
Transformer FeatO,
Attention
Transformer ReinL IU,
MIMIC
BLEU,
R-L,
METEOR,
CE
Wang et al. [2024b] Chest X-ray,
Chest CT
ResNet+
Transformer
AT Transformer ReLoss IU,
COV
BLEU,
R-L,
METEOR,
Grading
Xue et al. [2024] Chest X-ray,
Term
- ResNet+
Transformer
- Transformer,
Attention
FeatO,
Attention
Transformer - IU,
MIMIC
BLEU,
R-L,
METEOR
Wang et al. [2024a] Chest X-ray Resizing,
Cropping
DenseNet+
Transformer
AT-Class R2Gen AT IU,
MIMIC
BLEU,
R-L,
METEOR,
C-D,
CE
Jin et al. [2024] Chest X-ray,
RealR
ResNet AT-Class Transformer FeatO,
Attention
Transformer IU,
MIMIC
BLEU,
R-L,
METEOR,
CE
Table A2: Comparisons of the model performance on the MIMIC-CXR Dataset. B1, B2, B3, B4, R-L, C-D, P, R, and F represent BLEU-1, BLEU-2, BLEU-3, BLEU-4, ROUGE-L, CIDEr-D, precision, recall, and F1 score, respectively. The best and second best results are highlighted. All values were extracted from their papers.
[…]
Findings Section
Chen et al. [2021] 0.353 0.218 0.148 0.106 0.278 0.142 - 0.334 0.275 0.278
Pino et al. [2021] - - - - 0.185 - 0.238 0.381 0.531 0.428
Song et al. [2022] 0.360 0.227 0.156 0.117 0.287 0.148 - 0.444 0.297 0.356
Impression + Findings Section
Wu et al. [2022] 0.340 0.212 0.145 0.103 0.270 0.139 0.109 - - -
Wang et al. [2022d] 0.351 0.223 0.157 0.118 0.287 - 0.281 - - -
Jia et al. [2022] 0.363 0.228 0.156 0.130 0.300 - - - - -
Unspecified generated sections
Liu et al. [2021a] 0.344 0.217 0.140 0.097 0.218 0.133 - - - -
Liu et al. [2021c] 0.350 0.219 0.152 0.109 0.283 0.151 - 0.352 0.298 0.303
Liu et al. [2021b] 0.360 0.224 0.149 0.106 0.284 0.149 - - - -
Li et al. [2023c] 0.360 0.231 0.162 0.119 0.298 0.153 0.217 - - -
Lin et al. [2023] 0.362 0.227 0.155 0.113 0.283 0.142 - - - -
Zhang et al. [2023b] 0.362 0.229 0.157 0.113 0.284 0.153 - 0.380 0.342 0.335
Yang et al. [2022] 0.363 0.228 0.156 0.115 0.284 - 0.203 0.458 0.348 0.371
Wang et al. [2023c] 0.363 0.235 0.164 0.118 0.301 0.136 - - - -
Liu et al. [2021d] 0.369 0.231 0.156 0.118 0.295 0.153 - 0.389 0.362 0.355
Xue et al. [2024] 0.372 0.233 0.154 0.112 0.286 0.152 - - - -
Wang et al. [2024a] 0.374 0.230 0.155 0.112 0.279 0.145 0.161 0.483 0.323 0.387
Zhang et al. [2023a] 0.376 0.233 0.157 0.113 0.276 0.144 - - - -
You et al. [2021] 0.378 0.235 0.156 0.112 0.283 0.158 - - - -
Qin and Song [2022] 0.381 0.232 0.155 0.109 0.287 0.151 - 0.342 0.294 0.292
Wu et al. [2023] 0.383 0.224 0.146 0.104 0.280 0.147 - - - 0.758
Yang et al. [2023] 0.386 0.237 0.157 0.111 0.274 - 0.111 0.420 0.339 0.352
Wang et al. [2023b] - - - 0.119 0.286 0.158 0.259 - - -
Wang et al. [2023d] 0.386 0.250 0.169 0.124 0.291 0.152 0.362 0.364 0.309 0.311
Liu et al. [2023b] 0.391 0.249 0.172 0.125 0.304 0.160 - - - -
Huang et al. [2023] 0.393 0.243 0.159 0.113 0.285 0.160 - 0.371 0.318 0.321
Wang et al. [2022b] 0.395 0.253 0.170 0.121 0.284 0.147 - - - -
Jin et al. [2024] 0.398 - - 0.112 0.268 0.157 - 0.501 0.509 0.476
Wang et al. [2022e] 0.413 0.266 0.186 0.136 0.298 0.170 0.429 - - -
Kong et al. [2022] 0.423 0.261 0.171 0.116 0.286 0.168 - 0.482 0.563 0.519